I need to find the solution set by hand and having difficulty;
3x^2+3y^2=27
3x^2+2y^2=23

Answer by Alan3354(69443)

3x^2+3y^2=27
3x^2+2y^2=23
------------------ Subtract
y^2 = 4
y = -2, +2
============
3x^2+3y^2=27
x^2+y^2 = 9
x = -sqrt(5), +sqrt(5)